OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

Try writing the equation in terms of x.

OpenStudy (anonymous):

can i just please have the answer..... :/ im really really frustrated... @NeetziD

OpenStudy (anonymous):

okay. But I prefer giving detail, because I can solve a problem or two, but during exams, you're on your own. The answer is -5/2

OpenStudy (anonymous):

these are just so i can study these are for a study guide so i can study them. im sorry for bothering you :( @NeetziD

OpenStudy (anonymous):

No, it's not bothering. Explaining stuff to you strengthens my concepts.

OpenStudy (anonymous):

Just it's up to you to follow and understand, or to settle for a solution.

OpenStudy (anonymous):

5y=-2x-10 y=-2/5x-2 slope=-2/5

OpenStudy (anonymous):

I gave slope wrt abscissa. :\

OpenStudy (anonymous):

@makamuniko gave it w.r.t ordinate.